# Semantic Depth Matters: Explaining Errors of Deep Vision Networks through Perceived Class Similarities

### v1

@@ 1 Introduction
Deep neural networks have revolutionized countless fields in computer vision. However, a significant challenge still lies in understanding the rationale for their decision-making. At the same time, this understanding is crucial for reliability and transparency given the growing responsibility delegated to the intelligent systems. While much effort is put to create more accurate systems and explain their single decisions [39], more emphasis should be placed on network errors. It is because it is not the correct predictions that make the network’s decisions intuitive, but precisely their mistakes.
There is increasing demand for methods with built-in interpretability alongside accuracy metrics. Cognitive psychology offers relevant concepts, such as similarities and hierarchies, which are related to categorization [28]. Although initially explored in network evaluation, research shifted primarily towards accuracy-driven performance improvements [29, 1]. We argue these concepts merit renewed attention, as simple metrics (accuracy, loss) provide limited insights into how networks perceive class similarities and their impact on decisions and errors – a shortcoming highlighted by recent studies [11, 2, 1]. While existing work acknowledges that models frequently confuse similar classes [21, 7], the degree to which perceived similarity affects errors and their predictability remains underexplored.
@@ 2 Related Work
Similarity and semantic hierarchies are vital tools in cognitive sciences to examine human concept representations [8, 28]. Similarity is considered in vertical and horizontal dimensions [28]: vertical captures categorical hierarchy, whereas horizontal reflects pairwise category relationships. The vertical dimension thus effectively assesses semantic understanding depth in deep networks [1]. Semantic similarity, often relating to stable lexical structures like WordNet [20], serves as a standard reference when comparing with neural representations [7, 2, 1]. However, alternatives like human ratings are also employed [23]. Network-derived similarities typically use stimulus-based methods involving extracted features or confusion matrices [11, 12, 13, 36]. Recently, image-free methods leveraging classifier weights have gained traction due to computational efficiency, particularly for adversarial evaluation metrics [6, 21]. Representing inter-class similarities as affinities within similarity graphs provides structured insight into concept relationships. Graph-based methods, widely used for analyzing internal neural connectivity and performance prediction [38], are less common in examining high-level concepts for explainability, with few exceptions utilizing stimulus-based representations [26]. Clustering and community detection effectively identify category relationships within graphs [15], but their use in explainability remains limited. Existing metrics mainly emphasize horizontal similarities (pairwise relations) [21, 6, 2], with fewer addressing hierarchical depth through common ancestors for assessing dataset difficulty [4] or severity of mistakes [1]. Visual examination commonly targets image-level region activations [40, 31], with limited studies addressing whole-category representations [2].
@@ 3.2 Class Similarity Graphs
Class Similarity Graphs (CSG) provide a representation of relationships between the known classes. Given a large set of concepts $\mathcal{Z}$, a set of classes $c\in\mathcal{C}\subset\mathcal{Z}$ can be represented by nodes of undirected, weighted graphs $\mathcal{G}=\langle\mathcal{C},\mathcal{E},\boldsymbol{S}\rangle$, where the non-negative, symmetric Class Similarity Matrices (CSM) $\boldsymbol{S}\in\mathbb{R}_{+}^{\mathcal{C}\times\mathcal{C}}$ unambiguously define sets of edges storing all pairwise similarities between classes =$\mathcal{E}=\{e:e\in\mathcal{C}\times\mathcal{C}\land\boldsymbol{S}(e)>0\}$, where $\boldsymbol{S}(e)$ denotes weight of a weight matrix $\boldsymbol{S}$ for an edge $e$ (note that this matrix is symmetric). Initially, we assume that these graphs are fully connected, but they can be sparsified as zero values in CSM implicate that the corresponding edge does not exist.
@@
  • Semantic Similarity Graph (SSG): based on semantic similarity computed using the WordNet [20] ontology-based distances (we use Path length [25], Wu-Palmer - WUP [37] and Leacock-Chodorow [16] - LCH).
  • Network Structural Similarity Graph (NSSG): Derived without input data, from the weights of the neural network. To obtain weight matrices, we use a final classification layer of a deep network, which weights can be treated as learned class representations [21]. To do this, each neuron of the classification layer is assumed to be corresponding to one of the classes $c\in\mathcal{C}$. A vector of weights $\boldsymbol{w}_{c}\in\mathbb{R}^{|\mathcal{C}|}$ connecting $c$-th neuron to the penultimate layer is treated as a representation of $c$ in the last layer’s feature space. Similarity between these ‘templates’ is computed with cosine similarity. Computing the similarities between all class pairs and scaling them to range $\langle 0,1\rangle$ results in the Structural CSM.
@@ 3.3 Similarity Depth (SD)
WordNet organizes nouns in a hypernym–hyponym hierarchy (ISA hierarchy) ranging from very general concepts (e.g., ‘entity’) at the top to highly specific ones (e.g., Irish wolfhound’) at the bottom [20]. It is generally assumed that synsets residing deeper in the WordNet hierarchy are semantically closer [35]. Consequently, depth within the ISA hierarchy serves as an indicator of both semantic similarity and concept specificity, underpinning various similarity metrics such as Wu and Palmer similarity [37].
@@ 3.4 Visual Similarity Graph Explanation
Next, we use the previously defined weights of the final classifier as class templates as the mean feature representations, $\boldsymbol{w}_{i}$ and $\boldsymbol{w}_{j}$ for classes $c_{i}$ and $c_{j}$ respectively. To determine the most and least supporting images (images from class $c_{i}$ supporting class $c_{j}$ and vice versa), we compute the similarity between each instance and the weight template of the opposite class using a cosine similarity, and for each image $x_{i,k}$ obtain similarity $s_{i,k}^{(j)}$ from class $c_{i}$, and for each image $l$ from class $c_{j}$ $s_{j,l}^{(i)}$. We then select:
@@ 3.5 Measuring the Compliance of Structure-based and Confusion-based Graphs
Similarity Graph Compliance (SGC) is a metric that measures the alignment between two similarity graphs, in particular, between similarity graphs of a given network created based on its internal representations’ similarity or similarity estimated based on its classification errors. High alignment between these two similarity graphs means that the network’s explicit internal similarity representation (direct estimation of similarity) matches its implicit judgments revealed by its mistakes (indirect estimation), which mirrors human cognitive consistency [8]. Since similarity representation can be derived directly from weights, if these two estimates align well, the network’s errors can be inferred from its internal structure without experiments and data. This allows the evaluation of an aligned network in an ad-hoc manner directly from a trained model. High alignment also suggests that the network’s internal class representations capture the key features of the underlying task performance and error patterns. In contrast, low alignment indicates that the model’s internal representations do not adequately reflect the reasons for its mistakes.
@@ 3.6 Experimental Procedure
Convolutional Neural Networks (CNNs) Mobilenet V2 [30], ResNet 18 [9],
DenseNet 121 [10], EfficientNet [32],
ConvNeXt [18]
Vision Transformers (ViTs) ViT B [5], SWIN T [17]
Hybrid model MaxVit [33]
Our training procedure (excluding the learning rate for transformer models) remained consistent to ensure reproducibility. We trained the models for 300 epochs with a batch size of 128. For optimization, we used AdamW optimizer [19] and added additional learning rate chained schedulers. For the warmup period, we used a linear scheduler for 10 epochs, and then Reduce on Plateau [22] scheduler with a 20-epoch patience period and a reduce factor of 0.8.
@@
In our experiments with different training paradigms than the supervised one, we use DINOv2 [24] trained with a Self-Supervised procedure and an image-text model - CLIP [27]. We take the pre-trained models from: DINOv2: https://huggingface.co/docs/transformers/model_doc/dinov2 and CLIP: https://huggingface.co/docs/transformers/model_doc/clip.
@@
The main body of our experiments was performed on two commonly used image classification datasets – Mini-ImageNet [34] and CIFAR-100 [14]. Mini-ImageNet dataset consists of 50000 training images and 10000 testing images, evenly distributed across 100 classes. It includes groups of objects in categories such as animals (e.g., tiger, whale), man-made objects (e.g. dome, tank), plants, and food entities. CIFAR-100 consists of 60000 32x32 px color images divided evenly into 100 classes, with entity labels similar to Mini-ImageNet. There are 600 images per class, with 500 images for training and 100 for testing.
@@
For computing semantic similarities with the following distance measures: path length (path) [25], Wu and Palmer (WUP) [37], Leacock and Chodorow (LCH) [16], and thus obtaining the Class Similarity Matrices, we use the NLTK framework [3] along with WordNet linguistic taxonomy. We compute semantic similarity for each pair of classes in the used datasets.
@@ 4 Results
Neural networks demonstrated lower semantic depths (approximately 7-8). To contextualize these values, Table 3 provides illustrative examples from mini-ImageNet. Semantic depth values 7-8 convey detailed and specific semantic information, whereas shallow depths indicate meaningless or excessively general similarities. This aligns closely with the observations reported in [35]. Its authors noted that for shallow LCS depths (less than 5), concept similarity is independent of depth. However, for depths beyond 5, common ancestors correlate with increased concept similarity. This supports our findings that neural networks tested in this study, on average, operate at a semantic depth indicative of specific semantic similarities likely correlating with human perception.
@@
mini-ImageNet WordNet Path 8.96 (15) 8.96 (15) 8.90 (20) 8.39 (4) 8.52 (1)
WordNet WUP 8.96 (15) 8.57 (6) 9.05 (14) 8.45 (4) 8.56 (1)
@@
DenseNet121 7.36 (10) 7.13 (1) 6.66 (14) 7.29 (3) 7.35 (1)
EfficientNetV2 7.43 (16) 7.08 (1) 7.12 (17) 6.92 (1) 7.18 (1)
MobileNetV2 7.32 (6) 7.34 (1) 7.47 (15) 7.36 (3) 7.62 (1)
ConvNeXt 7.43 (8) 7.33 (1) 7.29 (15) 7.20 (3) 7.34 (1)
ResNet152 7.70 (8) 7.64 (1) 7.41 (18) 7.44 (3) 7.64 (1)
MaxVitT 7.43 (16) 7.08 (1) 7.12 (17) 7.08 (1) 7.18 (1)
Swin-base 7.57 (8) 7.29 (1) 7.20 (15) 7.27 (3) 7.29 (1)

### v2

@@ 1 Introduction
Deep neural networks have revolutionized countless fields in computer vision. However, a significant challenge still lies in understanding the rationale for their decision-making. At the same time, this understanding is crucial for reliability and transparency given the growing responsibility delegated to the intelligent systems. While much effort is put to create more accurate systems and explain their single decisions [bib:yuan2021explaining], more emphasis should be placed on network errors. It is because it is not the correct predictions that make the network’s decisions intuitive, but precisely their mistakes.
There is increasing demand for methods with built-in interpretability alongside accuracy metrics. Cognitive psychology offers relevant concepts, such as similarities and hierarchies, which are related to categorization [bib:rosch1978cognition]. Although initially explored in network evaluation, research shifted primarily towards accuracy-driven performance improvements [bib:russakovsky2015imagenet, bib:bertinetto2020making]. We argue these concepts merit renewed attention, as simple metrics (accuracy, loss) provide limited insights into how networks perceive class similarities and their impact on decisions and errors – a shortcoming highlighted by recent studies [bib:huang2021semantic, bib:bilal2017convolutional, bib:bertinetto2020making]. While existing work acknowledges that models frequently confuse similar classes [bib:mopuri2020adversarial, bib:filus2024similarity], the degree to which perceived similarity affects errors and their predictability remains underexplored.
@@ 2 Related Work
Similarity and semantic hierarchies are vital tools in cognitive sciences to examine human concept representations [bib:goldstone1994similarity, bib:rosch1978cognition]. Similarity is considered in vertical and horizontal dimensions [bib:rosch1978cognition]: vertical captures categorical hierarchy, whereas horizontal reflects pairwise category relationships. The vertical dimension thus effectively assesses semantic understanding depth in deep networks [bib:bertinetto2020making]. Semantic similarity, often relating to stable lexical structures like WordNet [bib:miller1990introduction], serves as a standard reference when comparing with neural representations [bib:filus2024similarity, bib:bilal2017convolutional, bib:bertinetto2020making]. However, alternatives like human ratings are also employed [bib:muttenthaler2024improving]. Network-derived similarities typically use stimulus-based methods involving extracted features or confusion matrices [bib:huang2021semantic, bib:kornblith2019similarity, bib:kriegeskorte2008rdm, bib:williamsequivalence]. Recently, image-free methods leveraging classifier weights have gained traction due to computational efficiency, particularly for adversarial evaluation metrics [bib:filus2023netsat, bib:mopuri2020adversarial]. Representing inter-class similarities as affinities within similarity graphs provides structured insight into concept relationships. Graph-based methods, widely used for analyzing internal neural connectivity and performance prediction [bib:you2020graph], are less common in examining high-level concepts for explainability, with few exceptions utilizing stimulus-based representations [bib:purvine2023experimental]. Clustering and community detection effectively identify category relationships within graphs [bib:lawson2012population], but their use in explainability remains limited. Existing metrics mainly emphasize horizontal similarities (pairwise relations) [bib:mopuri2020adversarial, bib:filus2023netsat, bib:bilal2017convolutional], with fewer addressing hierarchical depth through common ancestors for assessing dataset difficulty [bib:deng2010does] or severity of mistakes [bib:bertinetto2020making]. Visual examination commonly targets image-level region activations [bib:zintgraf2017visualizing, bib:selvaraju2016grad], with limited studies addressing whole-category representations [bib:bilal2017convolutional].
@@ 3.2 Class Similarity Graphs
Class Similarity Graphs (CSG) provide a representation of relationships between the known classes. Given a large set of concepts $\mathcal{Z}$, a set of classes $c\in\mathcal{C}\subset\mathcal{Z}$ can be represented by nodes of undirected, weighted graphs $\mathcal{G}=\langle\mathcal{C},\mathcal{E},\boldsymbol{S}\rangle$, where the non-negative, symmetric Class Similarity Matrices (CSM) $\boldsymbol{S}\in\mathbb{R}_{+}^{\mathcal{C}\times\mathcal{C}}$ unambiguously define sets of edges storing all pairwise similarities between classes $\mathcal{E}=\{e:e\in\mathcal{C}\times\mathcal{C}\land\boldsymbol{S}(e)>0\}$, where $\boldsymbol{S}(e)$ denotes weight of a weight matrix $\boldsymbol{S}$ for an edge $e$ (note that this matrix is symmetric). Initially, we assume that these graphs are fully connected, but they can be sparsified as zero values in CSM implicate that the corresponding edge does not exist.
@@
  • Semantic Similarity Graph (SSG): based on semantic similarity computed using the WordNet [bib:miller1990introduction] ontology-based distances (we use Path length [bib:pedersen2004wordnet], Wu-Palmer - WUP [bib:wu:1994verb] and Leacock-Chodorow [bib:leacock1998combining] - LCH).
  • Network Structural Similarity Graph (NSSG): Derived without input data, from the weights of the neural network. To obtain weight matrices, we use a final classification layer of a deep network, which weights can be treated as learned class representations [bib:mopuri2020adversarial]. To do this, each neuron of the classification layer is assumed to be corresponding to one of the classes $c\in\mathcal{C}$. A vector of weights $\boldsymbol{w}_{c}\in\mathbb{R}^{|\mathcal{C}|}$ connecting $c$-th neuron to the penultimate layer is treated as a representation of $c$ in the last layer’s feature space. Similarity between these ‘templates’ is computed with cosine similarity. Computing the similarities between all class pairs and scaling them to range $\langle 0,1\rangle$ results in the Structural CSM.
@@ 3.3 Similarity Depth (SD)
WordNet organizes nouns in a hypernym–hyponym hierarchy (ISA hierarchy) ranging from very general concepts (e.g., ‘entity’) at the top to highly specific ones (e.g., Irish wolfhound’) at the bottom [bib:miller1990introduction]. It is generally assumed that synsets residing deeper in the WordNet hierarchy are semantically closer [bib:wang2011refining]. Consequently, depth within the ISA hierarchy serves as an indicator of both semantic similarity and concept specificity, underpinning various similarity metrics such as Wu and Palmer similarity [bib:wu:1994verb].
@@ 3.4 Visual Similarity Graph Explanation
Next, we use the previously defined weights of the final classifier as class templates as the mean feature representations, $\boldsymbol{w}_{i}$ and $\boldsymbol{w}_{j}$ for classes $c_{i}$ and $c_{j}$ respectively. To determine the most and least supporting images (images from class $c_{i}$ supporting class $c_{j}$ and vice versa), we compute the similarity between each instance and the weight template of the opposite class using a cosine similarity, and for each image $x_{i,k}$ obtain similarity $s_{i,k}^{(j)}$ from class $c_{i}$, and for each image $l$ from class $c_{j}$ - $s_{j,l}^{(i)}$. We then select:
@@ 3.5 Measuring the Compliance of Structure-based and Confusion-based Graphs
Similarity Graph Compliance (SGC) is a metric that measures the alignment between two similarity graphs, in particular, between similarity graphs of a given network created based on its internal representations’ similarity or similarity estimated based on its classification errors. High alignment between these two similarity graphs means that the network’s explicit internal similarity representation (direct estimation of similarity) matches its implicit judgments revealed by its mistakes (indirect estimation), which mirrors human cognitive consistency [bib:goldstone1994similarity]. Since similarity representation can be derived directly from weights, if these two estimates align well, the network’s errors can be inferred from its internal structure without experiments and data. This allows the evaluation of an aligned network in an ad-hoc manner directly from a trained model. High alignment also suggests that the network’s internal class representations capture the key features of the underlying task performance and error patterns. In contrast, low alignment indicates that the model’s internal representations do not adequately reflect the reasons for its mistakes.
@@ 3.6 Experimental Procedure
Convolutional Neural Networks (CNNs) Mobilenet V2 [mobilenet_model], ResNet 18 [resnet_model],
DenseNet 121 [densenet_model], EfficientNet [efficientnet_model],
ConvNeXt [convnext_model]
Vision Transformers (ViTs) ViT B [vit_model], SWIN T [swin_model]
Hybrid model MaxVit [maxvit_model]
Our training procedure (excluding the learning rate for transformer models) remained consistent to ensure reproducibility. We trained the models for 300 epochs with a batch size of 128. For optimization, we used AdamW optimizer [adamW] and added additional learning rate chained schedulers. For the warmup period, we used a linear scheduler for 10 epochs, and then Reduce on Plateau [rop_scheduler] scheduler with a 20-epoch patience period and a reduce factor of 0.8.
@@
In our experiments with different training paradigms than the supervised one, we use DINOv2 [bib:oquab2023dinov2] trained with a Self-Supervised procedure and an image-text model - CLIP [bib:radford2021learning]. We take the pre-trained models from: DINOv2: https://huggingface.co/docs/transformers/model_doc/dinov2 and CLIP: https://huggingface.co/docs/transformers/model_doc/clip.
@@
The main body of our experiments was performed on two commonly used image classification datasets – Mini-ImageNet [mini_imagenet] and CIFAR-100 [cifar]. Mini-ImageNet dataset consists of 50000 training images and 10000 testing images, evenly distributed across 100 classes. It includes groups of objects in categories such as animals (e.g., tiger, whale), man-made objects (e.g. dome, tank), plants, and food entities. CIFAR-100 consists of 60000 32x32 px color images divided evenly into 100 classes, with entity labels similar to Mini-ImageNet. There are 600 images per class, with 500 images for training and 100 for testing.
@@
For computing semantic similarities with the following distance measures: path length (path) [bib:pedersen2004wordnet], Wu and Palmer (WUP) [bib:wu:1994verb], Leacock and Chodorow (LCH) [bib:leacock1998combining], and thus obtaining the Class Similarity Matrices, we use the NLTK framework [bird2009natural] along with WordNet linguistic taxonomy. We compute semantic similarity for each pair of classes in the used datasets.
@@ 4 Results
Neural networks demonstrated lower semantic depths (approximately 7-8). To contextualize these values, Table 3 provides illustrative examples from mini-ImageNet. Semantic depth values 7-8 convey detailed and specific semantic information, whereas shallow depths indicate meaningless or excessively general similarities. This aligns closely with the observations reported in [bib:wang2011refining]. Its authors noted that for shallow LCS depths (less than 5), concept similarity is independent of depth. However, for depths beyond 5, common ancestors correlate with increased concept similarity. This supports our findings that neural networks tested in this study, on average, operate at a semantic depth indicative of specific semantic similarities likely correlating with human perception.
@@
mini-ImageNet WordNet Path 8.96 (15) 8.48 (2) 8.90 (20) 8.39 (4) 8.52 (1)
WordNet WUP 8.65 (3) 8.57 (5) 9.05 (14) 8.56 (4) 8.56 (1)
@@
DenseNet121 7.36 (10) 7.13 (1) 6.66 (14) 7.20 (3) 7.13 (1)
EfficientNetV2 7.61 (6) 7.56 (2) 7.68 (16) 7.34 (3) 7.62 (1)
MobileNetV2 7.32 (6) 7.34 (1) 7.47 (15) 7.36 (3) 7.34 (1)
ConvNeXt 7.51 (8) 7.33 (1) 7.29 (15) 7.33 (1) 7.33 (1)
ResNet152 7.57 (9) 7.29 (1) 7.20 (15) 7.27 (3) 7.29 (1)
MaxVitT 7.43 (16) 7.08 (1) 7.12 (17) 6.68 (3) 7.08 (1)
Swin-base 7.65 (7) 7.58 (1) 7.66 (18) 7.39 (4) 7.58 (1)